# Position: Beyond Assistance – Reimagining LLMs as Ethical and Adaptive Co-Creators in Mental Health Care

### v1

@@ Abstract
This position paper argues for a fundamental shift in how Large Language Models (LLMs) are integrated into the mental health care domain. We advocate for their role as co-creators rather than mere assistive tools. While LLMs have the potential to enhance accessibility, personalization, and crisis intervention, their adoption remains limited due to concerns about bias, evaluation, over-reliance, dehumanization, and regulatory uncertainties. To address these challenges, we propose two structured pathways: SAFE-i (Supportive, Adaptive, Fair, and Ethical Implementation) Guidelines for ethical and responsible deployment, and HAAS-e (Human-AI Alignment and Safety Evaluation) Framework for multidimensional, human-centered assessment. SAFE-i provides a blueprint for data governance, adaptive model engineering, and real-world integration, ensuring LLMs align with clinical and ethical standards. HAAS-e introduces evaluation metrics that go beyond technical accuracy to measure trustworthiness, empathy, cultural sensitivity, and actionability. We call for the adoption of these structured approaches to establish a responsible and scalable model for LLM-driven mental health support, ensuring that AI complements—rather than replaces—human expertise.
@@ 1 Introduction
The under-utilization of AI in mental health is not merely a technological issue but a reflection of deeper concerns surrounding trust, ethical considerations, and the preservation of human expertise (Hamdoun et al., 2023). As LLMs become increasingly sophisticated, the mental health community faces a critical challenge: how to leverage their transformative potential while upholding the human-centered principles that define effective care (Obradovich et al., 2024). This tension is further exacerbated by the ability of LLMs to mimic human interaction and generate seemingly personalized responses, which may lead individuals to overestimate the depth of understanding these models possess (Sharma et al., 2020). Such dynamics can result in undue trust in LLM outputs, potentially neglecting other forms of support or treatment (Hua et al., 2024b). These concerns are shared by both individuals seeking mental health support and the professionals providing it, creating resistance and uncertainty around AI integration (Sobaih et al., 2025). Without a clear framework to ensure complementarity between AI and human-led interventions, these technologies risk being underutilized or misapplied, undermining their potential to augment mental health. Despite these challenges, early applications of human-AI collaboration demonstrate promising results. For instance, HAILEY (Sharma et al., 2023), a system designed to enhance empathy in peer-to-peer mental health support, has shown that conversations co-authored by AI are consistently rated as more empathic and supportive than human-only interactions.
@@
However, the deployment of LLMs in mental health care remains fraught with technical and ethical challenges. Studies reveal that these models often exhibit demographic biases, producing less empathetic or even harmful responses when interacting with underrepresented groups (Zack et al., 2024; Raza et al., 2024). Furthermore, proprietary models, such as ChatGPT 3.5, have demonstrated unsafe triage rates, misclassify urgent mental health crises, and potentially delay critical care—raising serious concerns about their reliability in high-stakes scenarios (Fraser et al., 2023). The absence of robust frameworks for development, evaluation, and deployment makes it difficult to ensure the effectiveness and safety of these tools. Accordingly, this paper proposes a path forward, redefining the role of LLMs in this sensitive domain through collaborative, ethical, and adaptive AI–human partnerships.
@@
This paper argues that LLMs have reached a pivotal stage where their implementation and evaluation of mental care is both viable and necessary. We advocate for reimagining LLMs as active co-creators rather than passive assistants, emphasizing supportive, collaborative, ethical, and adaptive AI-human partnerships that enhance - rather than replace - human-led mental health support.
@@
  • Introducing the HAAS-e ((Human-AI Alignment and Safety Evaluation) Framework to rigorously assess LLMs in mental health using a multidimensional approach. It defines four core evaluation criteria—trustworthiness, fairness, empathy, and helpfulness—operationalized through four novel quantitative metrics that measure alignment with human expertise, cultural sensitivity, personalization effectiveness, and actionability. Additionally, it integrates four validation methods—randomized trials, multi-method evaluations, red teaming, and adversarial testing—to ensure safety, accountability, and real-world applicability as shown in Figure 1.
@@
2 Alternative Views on LLMs in Mental Health
@@
Response: To mitigate over-reliance, it is essential to implement LLMs as complementary tools rather than replacements for human therapists (Sharma et al., 2023). Educating users on limitations, personalizing care strategies, and integrating feedback mechanisms ensure adaptation to individual needs while also encouraging users to seek human support when necessary (Strong et al., 2024). LLM systems can provide initial support when we integrate safety nets and escalation protocols while ensuring high-risk cases are promptly addressed by qualified professionals.
@@ 3 Prior Efforts in LLM-Powered Applications for Mental Health: A Landscape
The growing demand for mental health services, exacerbated by the COVID-19 pandemic (Hamdoun et al., 2023), has led to the exploration of generative AI technologies in various mental health applications (Zhang & Boulos, 2023; C. & John, 2023). One of the core technologies used in the Generative AI domain is LLMs, such as ChatGPT and GPT-4 (OpenAI, 2023), which utilize billions of parameters to generate coherent, contextually appropriate responses in mental health dialogues (Guo et al., 2024b; Torous & Blease, 2024). LLMs have been effectively applied in various application areas of mental health, such as crisis intervention (Obadinma et al., 2024; Sharma et al., 2024), therapy recommendations (Wilhelm et al., 2023; M. & Daniel, 2024; Berrezueta-Guzman et al., 2024), etc. Other applications of LLMs in mental healthcare include the work of Perlis et al. (2024), where they showed GPT-4 aligns with expert bipolar depression management, while  Lee et al. (2024) found GPT-4 had comparable sensitivity to clinicians in predicting suicidal ideation from intake data. Moreover, domain-specific LLMs have also gained a lot of attention recently in the mental healthcare domain (Yang et al., 2024). For instance, the Serena model, (L. et al., 2023) is developed as an effective counselor and demonstrates enhanced relevance and sensitivity toward therapeutic approaches (L. et al., 2023) with just 2.7 billion-parameters. More recently, Guo et al. (2024a) introduced SouLLMate, an adaptive LLM system integrating Retrieval-Augmented Generation (RAG), suicide risk detection, and proactive dialogues to enhance accessibility in mental health support.
In mental health applications, conversational AI tasks represent a major application area, with chatbots designed to engage users in text-based therapeutic conversations or monitor mental well-being (Liu et al., 2024a). For example, the chatbot Woebot, which uses cognitive-behavioral techniques, has shown efficacy in alleviating symptoms of depression and anxiety by delivering timely interventions (Fitzpatrick et al., 2017). The SuDoSys chatbot (Chen et al., 2024), which is based on WHO’s PM+ framework, ensures structured multi-turn psychological counseling with coherent stage tracking. The Coral framework proposed by Harsh et al. (2021) is designed to integrate conversational agents in mental health applications. For the evaluation of LLMs in clinical conversations, Johri et al. (2025) present CRAFT-MD, an evaluation framework assessing diagnostic reasoning in clinical LLMs, highlighting limitations of LLMs in conversational accuracy and the need for multimodal integration before deployment.
@@ 4 Key Challenges in Utilizing LLMs for Mental Health
LLMs in mental health applications face significant risks related to correctness, safety, and therapeutic reliability. Issues such as hallucinations, misinformation, and inappropriate responses (Zhao et al., 2023) necessitate more structured engineering processes (e.g., construction of optimized prompts) and real-world testing to ensure reliability and alignment with mental health best practices. Researchers emphasize the need for careful implementation, collaboration with stakeholders, and integration into existing healthcare systems (Torous & Blease, 2024). As the field evolves, there is a focus on developing empathetic, context-aware conversational agents (Harsh et al., 2021) and exploring diverse applications of AI in healthcare (Gozalo-Brizuela & Garrido-Merchan, 2023).
Moreover, ChatGPT-like closed-source proprietary LLMs are only accessible via their APIs (Laskar et al., 2024), which restricts users from fine-tuning the models locally or accessing their internal layers and weights (Pfohl et al., 2024). Moreover, relying too much on APIs raises privacy and security concerns, as sensitive data must be shared with third-party providers, increasing risks of data exposure. The lack of transparency in these models further complicates efforts to thoroughly evaluate their reliability and safety, a critical issue when handling sensitive mental health information (Lawrence et al., 2024).
@@
There is also a growing imbalance in AI accessibility across different demographics and healthcare systems. For instance, the cost of fine-tuning models for specific populations remains prohibitively high, leading to disparities in how well these tools serve different groups (Obradovich et al., 2024). While a recent study by  Stade et al. (2024) proposed a framework for the responsible development of LLMs in behavioral healthcare that could potentially augment or even replace certain aspects of human-led psychotherapy, the authors also acknowledge significant ethical and practical challenges associated with implementing this framework. Additionally, the over-alignment of models to safety constraints can result in over-cautious responses, where LLMs refuse to engage with critical mental health queries, limiting their usefulness in real therapeutic settings (Obradovich et al., 2024).
@@ 5.1 The Data Foundation: Preparing Reliable and Diverse Mental Health Data
A core component is population variability, where the source data should reflect different demographics, cultural backgrounds, languages, and mental health conditions (Obermeyer et al., 2019). Regular audits must be conducted to identify the over-representation or under-representation of specific groups (Mienye et al., 2024). Adoptive sampling strategies (Lum & Johndrow, 2016; Chawla et al., 2002) or synthetic data augmentation (Shahul Hameed et al., 2024; Juwara et al., 2024) should be employed where necessary to correct disparities, ensuring mitigating the risk of biases and fairness across wide audience (Abràmoff et al., 2023; Zhao et al., 2024).
@@ 5.2 Model Engineering: Designing Adaptive and Effective LLMs
LLMs designed for mental health must be continuously specialized and refined to maintain therapeutic relevance, ethical integrity, and cultural competence (Guo et al., 2024b; Thakkar et al., 2024). Adopting (e.g., fine-tuning or instruction-tuning) high-quality and domain-specific datasets is essential to embed empathy, rapport-building, and risk assessment into model behavior (Yang et al., 2024). Expert-in-the-loop mechanisms must be integrated to ensure sustained alignment with real-world therapeutic practices, allowing for iterative refinement based on feedback and emerging patient needs (Guo et al., 2024a). Furthermore, models must dynamically adapt to linguistic evolution, cultural shifts, and emerging mental health concerns, ensuring that LLM remains an inclusive, context-aware, and reliable support tool (Stade et al., 2024; Thakkar et al., 2024).
@@
Effective mental health AI requires carefully designed prompts for model adoptations and tuning that shape interactions in a supportive and actionable manner (Li et al., 2024b; Yu & McGuinness, 2024; Priyadarshana et al., 2024). Empathy-driven prompts position the LLM as a compassionate ally, encouraging users to share their feelings safely. Scenario-specific templates address diverse mental health contexts, from anxiety management to crisis support. Prompts also include calls to action, encouraging users to take steps (e.g., contacting a helpline), making the system both informative and actionable (Meskó, 2023; Patil et al., 2024).
@@
Tree of Thoughts (ToT) (Yao et al., 2024) and Chain of Thought (CoT) (Wei et al., 2022) reasoning enhance AI ability to break down complex mental health queries into structured, transparent decision paths, improving logical coherence and reducing hallucinations (Yao et al., 2023). By guiding the model to think through problems systematically rather than relying on direct pattern matching, these techniques help reduce hallucinations and enhance interpretability. Moreover, research on self-reflective AI suggests that LLMs can improve their accuracy by critically evaluating their own outputs before finalizing responses (Ji et al., 2023; Shinn et al., 2024). Furthermore, integrating uncertainty-aware architectures further enhances safety by enabling models to quantify their confidence levels in sensitive conversations (Yin et al., 2024). When faced with high-risk inputs, these architectures allow AI systems to flag uncertain responses for human review, reducing the likelihood of misleading or inadequate crisis interventions.

### v2

@@ Abstract
This position paper argues for a fundamental shift in how Large Language Models (LLMs) are integrated into the mental health care domain. We advocate for their role as co-creators rather than mere assistive tools. While LLMs have the potential to enhance accessibility, personalization, and crisis intervention, their adoption remains limited due to concerns about bias, evaluation, over-reliance, dehumanization, and regulatory uncertainties. To address these challenges, we propose two structured pathways: SAFE-i (Supportive, Adaptive, Fair, and Ethical Implementation) Guidelines for ethical and responsible deployment, and HAAS-e (Human-AI Alignment and Safety Evaluation) Framework for multidimensional, human-centered assessment. SAFE-i provides a blueprint for data governance, adaptive model engineering, and real-world integration, ensuring LLMs align with clinical and ethical standards. HAAS-e introduces evaluation metrics that go beyond technical accuracy to measure trustworthiness, empathy, cultural sensitivity, and actionability. We call for the adoption of these structured approaches to establish a responsible and scalable model for LLM-driven mental health support, ensuring that AI complements, rather than replaces human expertise.
@@ 1 Introduction
The under-utilization of AI in mental health is not merely a technological issue but a reflection of deeper concerns surrounding trust, ethical considerations, and the preservation of human expertise (Hamdoun et al., 2023). As LLMs become increasingly sophisticated, the mental health community faces a critical challenge: how to leverage their transformative potential while upholding the human-centered principles that define effective care (Obradovich et al., 2024). This tension is further exacerbated by the ability of LLMs to mimic human interaction and generate seemingly personalized responses, which may lead individuals to overestimate the depth of understanding these models possess (Sharma et al., 2020). Such dynamics can result in undue trust in LLM outputs, potentially neglecting other forms of support or treatment (Hua et al., 2024b).
Furthermore, recent studies show increasing public trust and optimism. Varghese et al. (2024) found that 53% of respondents moderately trust AI tools for mental health, valuing their accessibility, anonymity, and stigma reduction, while 34.8% expressed optimism due to their constant availability and cost-effectiveness. Alanezi (2024) reported positive perceptions of ChatGPT for psychoeducation and emotional support, and Siddals et al. (2024) observed high user engagement and therapeutic benefits from generative AI chatbots. Together, these findings highlight AI’s growing acceptance, particularly when systems are empathetic, support appropriate crisis escalation, and clearly define AI-human boundaries. This supports the concept of user acceptability of the use of LLMs in mental health, which is crucial for the success of AI-driven mental health interventions.
Through our collaboration with an e-mental health organization, where we evaluated LLMs on anonymized crisis support conversations, we found that the lack of robust development, evaluation, and deployment frameworks with human-in-the-loop hinders their safe and effective integration in mental health care (Obadinma et al., 2025). These concerns are shared by both individuals seeking mental health support and the professionals providing it, creating resistance and uncertainty around AI integration (Sobaih et al., 2025). Without a clear framework to ensure complementarity between AI and human-led interventions, these technologies risk being underutilized or misapplied, undermining their potential to augment mental health. Despite these challenges, early applications of human-AI collaboration demonstrate promising results. For instance, HAILEY (Sharma et al., 2023), a system designed to enhance empathy in peer-to-peer mental health support, has shown that conversations co-authored by LLMs are consistently rated as more empathic and supportive than human-only interactions.
@@
However, the deployment of LLMs in mental health care remains fraught with technical and ethical challenges. Studies reveal that these models often exhibit demographic biases, producing less empathetic or even harmful responses when interacting with underrepresented groups (Zack et al., 2024; Raza et al., 2024). Furthermore, proprietary models, such as ChatGPT 3.5, have demonstrated unsafe triage rates, misclassify urgent mental health crises, and potentially delay critical care, raising concerns about their reliability in high-stakes scenarios (Fraser et al., 2023). The absence of robust frameworks for development, evaluation, and deployment makes it difficult to ensure the effectiveness and safety of these tools. Accordingly, this paper proposes a path forward, redefining the role of LLMs in this sensitive domain through collaborative, ethical, and adaptive AI–human partnerships.
@@
This paper argues that LLMs have reached a pivotal stage where their implementation and evaluation of mental care is both viable and necessary. We advocate for reimagining LLMs as active co-creators rather than passive assistants, emphasizing supportive, ethical, and adaptive AI-human partnerships that enhance - rather than replace - human-led mental health support.
@@
  • Introducing the HAAS-e (Human-AI Alignment and Safety Evaluation) Framework to rigorously assess LLMs in mental health using a multidimensional approach. It defines four core evaluation criteria, including trustworthiness, fairness, empathy, and helpfulness, operationalized through four novel quantitative metrics that measure alignment with human expertise, cultural sensitivity, personalization, and actionability. Additionally, it integrates four validation methods—randomized trials, multi-method evaluations, red teaming, and testing—to ensure safety, accountability, and real-world applicability as shown in Figure 1.
@@
2 Alternative Views
@@
Response: To mitigate over-reliance, it is essential to implement LLMs as complementary tools rather than replacements for human therapists (Sharma et al., 2023). Educating users on limitations, personalizing care strategies, and integrating feedback mechanisms ensure adaptation to individual needs and encourage users to seek human support when necessary (Strong et al., 2024). LLM systems can provide initial support when we integrate safety nets and escalation protocols while ensuring high-risk cases are promptly addressed by qualified professionals.
@@ 3 Prior Efforts in LLM-Powered Applications for Mental Health: A Landscape
The growing demand for mental health services, exacerbated by the COVID-19 pandemic (Hamdoun et al., 2023), has led to the exploration of generative AI technologies in various mental health applications (Zhang & Boulos, 2023; C. & John, 2023). One of the core technologies used in the Generative AI domain is LLMs, such as ChatGPT and GPT-4 (OpenAI, 2023), which utilize billions of parameters to generate coherent, contextually appropriate responses in mental health dialogues (Guo et al., 2024b; Torous & Blease, 2024). LLMs have been effectively applied in various application areas of mental health, such as crisis intervention (Obadinma et al., 2025; Sharma et al., 2024), therapy recommendations (Wilhelm et al., 2023; Malgaroli & McDuff, 2024; Berrezueta-Guzman et al., 2024), etc.
Other applications of LLMs in mental healthcare include the work of Perlis et al. (2024), where they showed GPT-4 aligns with expert bipolar depression management, while  Lee et al. (2024) found GPT-4 had comparable sensitivity to clinicians in predicting suicidal ideation from intake data. Moreover, domain-specific LLMs have also gained a lot of attention recently in the mental healthcare domain (Yang et al., 2024). For instance, the Serena model, (Brocki et al., 2023) is developed as an effective counselor and demonstrates enhanced relevance and sensitivity toward therapeutic approaches (Brocki et al., 2023) with just 2.7 billion-parameters. More recently, Guo et al. (2024a) introduced SouLLMate, an adaptive LLM system integrating Retrieval-Augmented Generation (Gao et al., 2023), suicide risk detection, and proactive dialogues to enhance accessibility in mental health support.
In mental health applications, conversational AI tasks represent a major application area, with chatbots designed to engage users in text-based therapeutic conversations or monitor mental well-being (Liu et al., 2024a). For example, the chatbot Woebot, which uses cognitive-behavioral techniques, has shown efficacy in alleviating symptoms of depression and anxiety by delivering timely interventions (Fitzpatrick et al., 2017). The SuDoSys chatbot (Chen et al., 2024), which is based on WHO’s PM+ framework, ensures structured multi-turn psychological counseling with coherent stage tracking. The Coral framework proposed by Sakhrani et al. (2021) is designed to integrate conversational agents in mental health applications. For the evaluation of LLMs in clinical conversations, Johri et al. (2025) present CRAFT-MD, an evaluation framework assessing diagnostic reasoning in clinical LLMs, highlighting limitations of LLMs in conversational accuracy and the need for multimodal integration before deployment.
@@ 4 Key Challenges in Utilizing LLMs for Mental Health
LLMs in mental health applications face significant risks related to correctness, safety, and therapeutic reliability. Issues such as hallucinations, misinformation, and inappropriate responses (Zhao et al., 2023) necessitate more structured engineering processes (e.g., construction of optimized prompts) and real-world testing to ensure reliability and alignment with mental health practices. Researchers emphasize the need for careful implementation, collaboration with stakeholders, and integration into existing healthcare systems (Torous & Blease, 2024). As the field evolves, there is a focus on developing empathetic, context-aware conversational agents (Sakhrani et al., 2021) and exploring diverse applications of AI in healthcare (Gozalo-Brizuela & Garrido-Merchan, 2023).
Moreover, ChatGPT-like closed-source proprietary LLMs are only accessible via their APIs (Laskar et al., 2023; Jahan et al., 2024), which restricts users from fine-tuning the models locally or accessing their internal layers and weights (Pfohl et al., 2024). Also, relying too much on APIs raises privacy and security concerns, as sensitive data must be shared with third-party providers, increasing risks of data exposure. The lack of transparency in these models further complicates efforts to thoroughly evaluate their reliability and safety, a critical issue when handling sensitive mental health information (Lawrence et al., 2024).
@@
There is also a growing imbalance in AI accessibility across different demographics and healthcare systems. For instance, the cost of fine-tuning models for specific populations remains prohibitively high, leading to disparities in how well these tools serve different groups (Obradovich et al., 2024). While a recent study by  Stade et al. (2024) proposed a framework for the responsible development of LLMs in behavioral healthcare that could potentially augment or even replace certain aspects of human-led psychotherapy, the authors also acknowledge significant ethical and practical challenges with implementing this framework.
Additionally, the over-alignment of models to safety constraints can result in over-cautious responses, where LLMs refuse to engage with critical mental health queries, limiting their usefulness in real therapeutic settings (Obradovich et al., 2024). Lastly, a fundamental risk with deploying LLM in mental health settings is hallucination, where models generate output with incorrect or misleading information (Huang et al., 2023). In mental health, this issue is particularly sensitive, as inaccurate guidance or misinformation can have immediate and severe consequences.
@@
Another concern in applying LLMs to mental health support is the potential for over-reliance on AI-driven interventions, which could inadvertently result in dehumanization or a reduction in meaningful human interactions (Zhang & Wang, 2024; Lawrence et al., 2024). Major impacts may include users trusting AI more than human counselors, emotional support provided by LLMs could be perceived as sufficient, and systemic overuse of AI in mental health could result in reduced funding or prioritization for human-led services.
@@ 5.1 The Data Foundation: Preparing Reliable and Diverse Mental Health Data
A core component is population variability, where the source data should reflect different demographics, cultural backgrounds, languages, and mental health conditions (Obermeyer et al., 2019). Regular audits must be conducted to identify the over-representation or under-representation of specific groups (Mienye et al., 2024). Adoptive sampling strategies (Lum & Johndrow, 2016; Chawla et al., 2002) or synthetic data augmentation (Shahul Hameed et al., 2024; Juwara et al., 2024) should be employed where necessary to correct disparities, ensuring the mitigation of the risk of biases and fairness across a wide audience (Abràmoff et al., 2023; Zhao et al., 2024).
@@ 5.2 Model Engineering: Designing Adaptive and Effective LLMs
LLMs designed for mental health must be continuously specialized and refined to maintain therapeutic relevance, ethical integrity, and cultural competence (Guo et al., 2024b; Thakkar et al., 2024). Adopting (e.g., fine-tuning or instruction-tuning) high-quality and domain-specific datasets is essential to embed empathy, rapport-building, and risk assessment into model behavior (Yang et al., 2024). Expert-in-the-loop mechanisms must be integrated to ensure sustained alignment with real-world therapeutic practices, allowing for iterative refinement based on feedback and emerging patient needs (Guo et al., 2024a). Furthermore, models must dynamically adapt to linguistic evolution, cultural shifts, age-specific informal expressions, and emerging mental health concerns, ensuring that LLM remains an inclusive, context-aware, and reliable support tool (Stade et al., 2024; Thakkar et al., 2024).
@@
Effective mental health AI requires carefully designed prompts for model adaptations and tuning that shape interactions in a supportive and actionable manner (Li et al., 2024b; Yu & McGuinness, 2024; Priyadarshana et al., 2024). Empathy-driven prompts position the LLM as a compassionate ally, encouraging users to share their feelings safely. Scenario-specific templates address diverse mental health contexts, from anxiety management to crisis support. Prompts also include calls to action, encouraging users to take steps (e.g., contacting a helpline), making the system both informative and actionable (Meskó, 2023; Patil et al., 2024).
@@
Tree of Thoughts (ToT) (Yao et al., 2024) and Chain of Thought (CoT) (Wei et al., 2022) reasoning enhance AI ability to break down complex mental health queries into structured, transparent decision paths, improving logical coherence and reducing hallucinations in emotionally sensitive contexts (Yao et al., 2024). By guiding the model to think through psychosocial or affective problems systematically rather than relying on direct pattern matching, these techniques help in critical therapeutic or crisis scenarios and enhance interpretability. Moreover, research on self-reflective AI suggests that LLMs can improve their accuracy by critically evaluating their own outputs before finalizing responses (Ji et al., 2023; Shinn et al., 2024). Furthermore, integrating uncertainty-aware architectures further enhances safety by enabling models to quantify their confidence levels in sensitive conversations (Yin et al., 2024). When faced with high-risk inputs, these architectures allow AI systems to flag uncertain responses for human review, reducing the likelihood of misleading or inadequate crisis interventions.